OpenStudy (neen9707):

Can someone show me how to find the slope for this thank you What is the slope of the line through the points (-3, -4) and (-1, 4)? m = _?_

OpenStudy (er.mohd.amir):

slope=(y2-y1)/(x2-x1) where (x1,y1) and (x2,y2) are points

OpenStudy (mathstudent55):

Subtract the two y-coordinates. Subtract the two x-coordinates in the same order. Divide the first difference by the second difference. That is the slope.

OpenStudy (minecraftaddictionz):

(\Color{Red}Text{Admin}\)

OpenStudy (mathstudent55):

\(\tt slope = \dfrac{difference~in~y}{difference~in~x}\)

OpenStudy (campbell_st):

ok.. so here is an easy visual method ... draw a simple number plane and plot the points  form a right triangle... the run is the distance from -3 to -1 the rise is the distance from -4 to 4 the the slope is \[m = \frac{rise}{run}\] hope it helps

OpenStudy (neen9707):

Ok so (-4) - 4 / (-1) - (-3)? and the result will be the answer?

OpenStudy (campbell_st):

no.. it won't result in the answer you need to check your information

OpenStudy (neen9707):

oh whoops i have to switch the y's around oops lol